OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\left(\begin{matrix}(1+x)^{7}-(1-x)^{7} \\ (1+x)^{5}-(1-x)^{5}\end{matrix}\right)\]

OpenStudy (anonymous):

a good guess will solve it immediately

OpenStudy (anonymous):

@satellite73 : lol and i need that guess! =.=

OpenStudy (turingtest):

is this supposed to be a fraction and not a matrix?

OpenStudy (turingtest):

\[\lim_{x\to0}{(1+x)^7-(1-x)^7\over(1+x)^5-(1-x)^5}\]?

OpenStudy (anonymous):

@TuringTest :yes it should be a fraction but i dont know how to make a fraction

OpenStudy (anonymous):

@TuringTest :yup thats it !

OpenStudy (anonymous):

(1^7-1^7)/(1^5-1^5) Not sure you're able to do it like this, just guessing! :D

OpenStudy (turingtest):

@Noliec no that's not how you do it, and that gives 0/0 please don't guess like that

OpenStudy (anonymous):

I guess you would have to expand all factors and solve it that way?

OpenStudy (turingtest):

no there is a better way I unfortunately feel like the answer is escaping me... embarrassing :(

OpenStudy (turingtest):

something to do with difference of squares I'm sure

OpenStudy (anonymous):

how about L'hopital rule..

OpenStudy (anonymous):

@TuringTest :I think that w should distribute the power ,like make the 7th power =(1-x)^2 (1-x)5

OpenStudy (anonymous):

7/5

OpenStudy (turingtest):

L'hospital will get very messy I think @Eyad but then what...? wolfram has its answer by expanding, so that is one way but there must be a better one

OpenStudy (turingtest):

yes that's the answer @cinar but how did you get it without the wolf?

OpenStudy (anonymous):

nope.. L'hopital

OpenStudy (turingtest):

ok let me try it I thought it would be messy....

OpenStudy (turingtest):

I still get 0/0 with l'hospital I must be missing something

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{7(1+x)^6+7(1-x)^6}{5(1+x)^4+5(1-x)^4}=7/5\]

OpenStudy (turingtest):

how did the negative turn positive o-0 ?

OpenStudy (turingtest):

oh nevermind

OpenStudy (anonymous):

because \[y= f(x)^n\] \[y'=n f(x)^{n-1}*f'(x)\]

OpenStudy (turingtest):

duh ok

OpenStudy (turingtest):

funny how wolf didn't use l'hospital usually it uses it every chance it gets oh well, thanks cinar!

OpenStudy (anonymous):

Thanks @cinar

OpenStudy (anonymous):

np